OpenStudy (anonymous):

write the equation in standard form & determine the center, & radius if it exists: 2x^2+2y^2+8x-12y-8=0

OpenStudy (anonymous):

=x^2+y^2+4x-6y-4=0 (x+2)^2 -4 +(y-3)^2 -9 -4=0 (x+2)^2 +(y-3)^2 =17 so, your radius: r=\[r=\sqrt{17}\] & center: (-2,3)